# Di-NeRF: Distributed NeRF for Collaborative Learning with Unknown Relative Poses

### v1

@@ III-D Distributed Optimization of NeRF
The optimization through ADMM alternates between variables, and C-ADMM allows the agents to share the intermediate optimized variables to achieve a consensus. The optimization process is composed of three steps, in each step, other variables are treated as constants (superscript (k) denotes the step number in updating the variable):
$\bullet\leavevmode\nobreak\ \theta_{i}$-minimization ($z_{ij}$ and $y_{i}$ are constants):
$\underset{T_{i}^{g(k)},\theta_{i}^{(k)}}{\operatorname{min}}\leavevmode\nobreak\ L_{i}(\theta_{i}^{(k)},T_{i}^{g(k)})+\theta_{i}^{(k)\top}y_{i}^{(k)}+\frac{\rho}{2}\sum_{j\in\mathcal{N}_{i}}\left\|\theta^{(k)}_{i}-z_{ij}^{(k)}\right\|_{2}^{2},$ (5)
$\bullet\leavevmode\nobreak\ z_{ij}$-minimization ($L_{i}$, $\theta_{i}$ and $y_{i}$ are constants):
$\underset{z_{ij}^{(k)}}{\operatorname{min}}{-z_{ij}^{(k)\top}y_{i}^{(k)}+\frac{\rho}{2}\sum_{j\in\mathcal{N}_{i}}\left\|\theta_{i}^{(k)}-z_{ij}^{(k)}\right\|_{2}^{2}},$ (6)
@@
$y_{i}^{(k)}\leftarrow y_{i}^{(k-1)}+\rho\sum_{j\in\mathcal{N}_{i}}{(\theta_{i}^{(k)}-z_{ij}^{(k)})},$ (7)
The optimization step alternates between minimizing the updated local objective function with respect to primal variables ($\theta_{i}$) and maximizing the updated local objective function with respect to the dual variable. The $z_{ij}$-minimization step can be simplified to the following update (see appendix -A):
Therefore the $\theta_{i}$-minimization becomes
$\underset{T_{i}^{g(k)},\theta_{i}^{(k)}}{\operatorname{min}}\leavevmode\nobreak\ L_{i}(\theta_{i}^{(k)},T_{i}^{g(k)})+\theta_{i}^{(k)\top}y_{i}^{(k)}+\frac{\rho}{2}\sum_{j\in\mathcal{N}_{i}}\left\|\theta^{(k)}_{i}-\bar{\theta}^{(k)}\right\|_{2}^{2}.$ (9)
Algorithm 1 summarizes the optimization procedures for local primal, dual variables, and relative poses for each robot. In line 1, The initial values for the NeRF parameters $\theta_{i}^{(0)}$ and the dual variable $y_{i}^{(0)}$ are both set to zero. The initial value for the relative poses - three rotation angles $R$ and three translation displacements $t$ are explained in Sec. III-B. The first robot undergoes the NeRF training process during the first iteration (approximately 200 steps (lines 4-11)). Subsequently, it shares the learned weights with other robots. The remaining robots engage in collaborative optimization, simultaneously refining NeRF weights and adjusting relative camera poses in a distributed training framework based on the robot $R^{g}$ weights (lines 12-19). This iterative process of communication and training persists until uniformity is achieved in both the weights and optimized relative camera poses across all robots. The relative pose between robot $i$ coordinate system and robot 1 coordinate system (global coordinate) $T_{i}^{g}\in SE(3)$ is optimized jointly in the distributed training (line 14).
@@
1:Initialization: $k\leftarrow 0,\theta_{i}^{(0)}\in\mathbb{R}^{n},y_{i}^{(0)}=0,T_{i}^{g}\in SE(3)$
@@
4:procedure robot $R^{g}$
@@
6:         $\theta_{i}^{(k+1)}=\underset{\theta_{i}}{\operatorname{argmin}}\left\{L_{i}\left(\theta_{i}\right)+\theta_{i}^{\top}y_{i}^{(k)}\right.$
7:         $+\rho\sum_{j\in\mathcal{N}_{i}}\left\|\theta_{i}-\frac{1}{2}\left(\theta_{i}^{(k)}+\theta_{j}^{(k)}\right)\right\|_{2}^{2}\}$
8:         Communicate $\mathcal{Q}_{i}^{(k)}$ to all $j$ in $\mathcal{N}_{i}$
9:         Receive $\mathcal{Q}_{j}^{(k)}$ from all $j$ in $\mathcal{N}_{i}$
10:         $y_{i}^{(k+1)}=y_{i}^{(k)}+\rho\sum_{j\in\mathcal{N}_{i}}(\theta_{i}^{(k+1)}-\theta_{j}^{(k+1)})$
11:         $k\leftarrow k+1$
12:while stopping condition not satisfied do
13:procedure Other Robots
14:     for $i$ in $\mathcal{V}$ do
15:         $\theta_{i}^{(k+1)},T_{i}^{g(k+1)}=\underset{\theta_{i},T_{i}^{g}}{\operatorname{argmin}}\left\{L_{i}\left(\theta_{i},T_{i}^{g}\right)+\theta_{i}^{\top}y_{i}^{(k)}\right.$
16:         $+\rho\sum_{j\in\mathcal{N}_{i}}\left\|\theta_{i}-\frac{1}{2}\left(\theta_{i}^{(k)}+\theta_{j}^{(k)}\right)\right\|_{2}^{2}\}$
17:         Communicate $\mathcal{Q}_{i}^{(k)}$ to all $j$ in $\mathcal{N}_{i}$
18:         Receive $\mathcal{Q}_{j}^{(k)}$ from all $j$ in $\mathcal{N}_{i}$
19:         $y_{i}^{(k+1)}=y_{i}^{(k)}+\rho\sum_{j\in\mathcal{N}_{i}}(\theta_{i}^{(k+1)}-\theta_{j}^{(k+1)})$
20:         $k\leftarrow k+1$
21:while stopping condition not satisfied do
@@ III-E Convergence Properties
The convergence of C-ADMM and its variants typically requires the dual variables’ sum to converge to zero, a condition challenging in unreliable networks. Moreover, the nonlinearity and nonconvexity of neural networks specially NeRF, preclude guaranteed global solutions and linear rates [11, 50]. Despite these issues, Algorithm 1 has proven effective in practice for distributed NeRF training, as shown in Sec. IV, converging to solutions equivalent to centralized methods.
@@ IV Experiments
This section presents the experimental results on synthetic and real-world datasets, including 1) comparing Di-NeRF and a centralized model based on several criteria, 2) examining the impact of the number of robots and communication graphs, 3) analyzing the impact of the amount of overlap between the robots’ trajectories on convergence, and 4) evaluation using Waymo Block-NeRF dataset [10]. The architecture of NeRF is based on the instant neural graphics primitives [25]. Technical details are described in Sec IV-E.
For the relative poses, the initial estimates for the three translational components were set to 50 cm, and the three rotational components were offset by 10 degrees from the ground truth. These represent the maximum deviations that Di-NeRF has been capable of rectifying.
IV-A Di-NeRF vs Centralized NeRF on Synthetic Dataset
@@
In this experiment, we use Di-NeRF to learn the RGB and density field of a three-dimensional environment in a distributed manner and compare the results with a centralized NeRF. The robots do not have access to a global coordinate frame, thus the relative poses are unknown. The datasets used are from the synthetic NeRF dataset [2], including Chair, Hotdog, and Lego sequences. We also use the long sequence of Barn from the real-world dataset of Tanks and Temples Benchmark [51]. For all the permutations, the configurations for centralized and Di-NeRF are the same, including the number of iterations (5000 steps). The number of images used in the centralized setup is around 100 to 150 for synthetic data and 300 for the Barn dataset. For multiple robots, the images are divided among the robots, in a controlled manner, such that the overlaps are controlled. In all experiments, we follow the official pre-processing procedures and train/test splits.
The metrics used to compare the performance of Di-NeRF with the centralized solution are the average rendering i) PSNRs, ii) SSIM, and iii) camera pose accuracy (in degrees and cm). The accuracy is determined by comparing the ground truth and the optimized pose values.
For the synthetic sequences, two robots are involved, and the dataset is split into two segments. For instance, in the case of the Chair sequence, robot $R^{1}$ exclusively observes the front of the chair. In contrast, robot $R^{2}$ focuses solely on the back (without any overlap in the robots’ trajectories). A similar pattern is applied to other synthetic sequences. The whole data are divided into two parts for synthetic dataset and with COLMAP the poses were estimated for each segment. Therefore, the coordinate system for each robot is different, and the relative pose is estimated by running Di-NeRF. In Fig. 3, the convergence of the relative camera pose for chair dataset can be seen.
@@
Figure 3: Optimizing the relative poses for two robots for chair from synthetic NeRF dataset. The relative translation and rotation estimates are plotted. The desired translation/rotation is marked by a dash line.
Qualitative results illustrating this setup can be found in Fig. 4. In Fig. 4-(first column), a sample of raw input images for two robots is shown. Since robot $R^{1}$ has no image from the back of the chair, it fails to reconstruct the back. The same is true for the robot $R^{2}$ reconstructing the front of the chair. This can be seen in Fig. 4-(middle column). Di-NeRF enables all robots to render the whole scene, see Fig. 4-(last column). Additionally, we present a comparison of PSNR and L2 loss in Fig. 5. The PSNR and L2 convergences for Di-NeRF and centralized training exhibit nearly identical behavior. Table I presents the quantitative results. In this table, the metrics separated by the /// sign show the values of the metric for each robot. Overall, our distributed joint optimization model achieves similar quality compared to the centralized one.
@@
Figure 4: Di-NeRF can reconstruct the entire scene via the collaboration of two robots, where each robot only sees a part of the scene. Individual training results in poor reconstruction quality for some areas of the scene, whereas Di-NeRF maintains good quality throughout. In each column, the left image is for robot $R^{1}$ and the right one is for robot $R^{2}$.
Figure 5: NeRF validation Loss and PSNR for chair dataset - Di-NeRF compared with centralized for two robots
@@
Table I: Image and relative camera pose metric performances for two robots, Lower/higher better, is indicated by $\downarrow$/$\uparrow$ respectively. PSNR and SSIM are reported for each robot. The relative pose accuracy is shown for $R^{2}$. $R^{1}$ is the global frame.
@@
PSNR $\uparrow$ SSIM $\uparrow$ $\delta R^{\circ}$ $\delta t$cm
Chair 31.25/31.31 33.02 0.922/0.926 0.952 0.70 1.02
Hotdog 29.40/29.32 32.63 0.924/0.914 0.963 0.61 0.93
Lego 28.15/28.17 29.67 0.929/0.935 0.941 0.55 0.89
Barn 28.13/28.13 29.94 0.879/0.869 0.894 0.63 0.95
IV-B Di-NeRF vs Centralized NeRF on Real-world Dataset
This experiment shows the performance of Di-NeRF not only on real-world data but also when the number of robots increases from two to five. This scenario applies to large-scale environments. Here, the Barn dataset from Tanks and Temples is used. There are 300 images in this sequence. Images are divided by the number of robots such that there is no common image in the local data of each robot, and the robot trajectories have zero overlaps. Fig. 6 shows the rendered images from each robot’s distributed trained NeRF. Each robot can render a novel view from any point. For instance, Robot $R^{3}$ has never seen the large bay doors but it is able to render an image when requested, shown in Fig. 6 (b).
Next, the number of robots in this experiment is altered. In Fig. 7, the trajectory for each robot is shown alongside the rendered images for each robot after collaborative training is completed. Each color represents the trajectory of a different robot. Table II presents the quantitative results. The rendering qualities and accuracy of relative camera pose in comparison to a centralized approach are evaluated. As anticipated with a distributed method, the outcomes closely resemble those achieved with the centralized approach.
@@ IV-B Di-NeRF vs Centralized NeRF on Real-world Dataset
Figure 6: In row (a), input images, sourced from the Tanks and Temple dataset (Barn) are shown, which is provided to each robot. Rows (b, c, d) showcase rendered images generated by Di-NeRF. In these rows, the robots collaboratively communicate to reconstruct the complete scene. Notably, the views depicted in rows b, c, and d are strategically chosen—each is only visible in the raw data of a specific robot but can be rendered by all robots collectively in the final reconstruction.
Table II: Image and relative camera pose metric performances for different numbers of robots for Barn (Tanks and Temples dataset), Lower/higher better, is indicated by $\downarrow$/$\uparrow$ respectively. PSNR and SSIM are reported for each robot. The relative pose optimization accuracy for all robots is shown, except for robot $R^{1}$ which is considered as the global coordinate.
Di-NeRF Centralized Di-NeRF Centralized Relative Camera Pose Error
No. Robots PSNR $\uparrow$ SSIM $\uparrow$ $\Delta$ rot (deg) $\downarrow$ $\Delta$ tran (cm) $\downarrow$
2 28.13/28.13 29.94 0.879/0.869 0.894 0.63 0.95
3 28.23/28.25/28.25 29.94 0.868/0.857/0.861 0.894 0.68/0.69 0.97/0.94
4 28.75/28.59/28.67/28.57 29.94 0.874/0.808/0.812/0.854 0.894 0.56/0.51/0.60 0.84/0.86/0.87
5 28.47/28.48/28.50/28.51/28.47 29.94 0.867/0.842/0.853/0.847/0.850 0.894 0.74/0.71/0.75/0.81 0.90/0.91/0.93/0.89
In Table III, we experiment with different connectivity graphs for robots. Five robots are used with the Barn sequence from Tanks and Temples. In the star connection, $R^{1}$ is in the center. In Table III, the average result for 5 robots is shown.
Table III: Mean validation result for Di-NeRF and different connectivity graphs. Barn from Tanks and Temples dataset and 5 robots are chosen in all connectivity types. Di-NeRF works well even if the network of robots is not fully connected.
Communication PSNR SSIM $\delta\leavevmode\nobreak\ R^{\circ}$downarrow $\delta\leavevmode\nobreak\ t$ (cm) $\downarrow$
Fully Connected 28.49 0.867 0.74 0.90
Ring 28.28 0.847 0.84 0.87
Star 28.31 0.859 0.75 0.88
Line 28.17 0.855 0.91 0.97
@@
Figure 7: Di-NeRF for different numbers of robots. All robots are fully connected, and the relative poses and NeRFs are trained jointly. On the left side of the image, the allocation of frames and poses from the dataset to the different robots are shown in different colors. None of the robots have common images, but in the end, all robots can render the whole scene.

### v2

@@ III-D Distributed Optimization of NeRF
The optimization through ADMM alternates between variables, and C-ADMM allows the agents to share the intermediate optimized variables to achieve a consensus. The variables are ${\Gamma}_{i}=\{\theta_{i},T_{i}^{g},\boldsymbol{\psi}_{i}\}$, $z_{ij}$, and $y_{i}$. Optimization of $z_{ij}$ has a closed-form solution, see the website of the project, (superscript (k) denotes the step number in updating the variable):
therefore, the optimization step alternates between minimizing the updated local objective function with respect to primal variables ( ${\Gamma}_{i}=\{\theta_{i},T_{i}^{g},\boldsymbol{\psi}_{i}\}$) and maximizing the updated local objective function with respect to the dual variable:
$\bullet\leavevmode\nobreak\ \theta_{i}$-minimization
[…]
@@
$y_{i}^{(k)}\leftarrow y_{i}^{(k-1)}+\rho\sum_{j\in\mathcal{N}_{i}}{(\theta_{i}^{(k)}-\bar{\theta}^{(k)})},$ (10)
Alg. 1 summarizes the optimization procedures for local primal, dual variables, and relative poses for each robot. In line 1, The initial values for the NeRF parameters $\theta_{i}^{(0)}$, the dual variable $y_{i}^{(0)}$, and the depth distortion parameters $\boldsymbol{\psi}_{i}^{(0)}$ are set to zero. The initial value for the relative poses - three rotation angles $R$ and three translation displacements $t$ are explained in Sec. IV. The first robot $R^{g}$ undergoes the NeRF training process (ln 6-7), communicates the learned weights with other robots (ln 10-11), and updates the dual variable (ln 12) for 200 steps. The remaining robots engage in collaborative optimization, simultaneously refining NeRF weights and adjusting relative camera poses and depth distortion parameters in a distributed training framework based on the weights of robot $R^{g}$ (lines 8-9 for NeRF training, lines 10-12 for communication and dual update). This iterative communication and training process persists until uniformity is achieved in both the weights and optimized relative camera poses across all robots. The relative pose between robot $i$’s coordinate system and robot $R^{g}$’s coordinate system $T_{i}^{g}\in SE(3)$ is optimized jointly in the distributed training (ln 9).
@@
1:Initialization: $k\leftarrow 0,\theta_{i}^{(0)}\in\mathbb{R}^{n},y_{i}^{(0)}=0,T_{i}^{g}\in SE(3)$, $\boldsymbol{\psi_{i}}\in\mathbb{R}^{m}$
@@
4:while stopping condition not satisfied do
@@
6:         if robot $i$ is $R^{g}$ then
7:              $\theta_{i}^{(k+1)}=\underset{\theta_{i}}{\operatorname{argmin}}\{L_{i}\left(\theta_{i}\right)+\theta_{i}^{\top}y_{i}^{(k)}$
8:           $+\rho\sum_{j\in\mathcal{N}_{i}}\left\|\theta_{i}-\bar{\theta}^{(k)}\right\|_{2}^{2}\}$
9:         else
[…]
12:         Communicate $\mathcal{Q}_{i}^{(k)}$ to all $j$ in $\mathcal{N}_{i}$
13:         Receive $\mathcal{Q}_{j}^{(k)}$ from all $j$ in $\mathcal{N}_{i}$
14:         $y_{i}^{(k+1)}=y_{i}^{(k)}+\rho\sum_{j\in\mathcal{N}_{i}}(\theta_{i}^{(k+1)}-\theta_{j}^{(k+1)})$
15:         $k\leftarrow k+1$
@@ III-E Convergence Properties
The convergence of C-ADMM methods typically requires the dual variables’ sum to converge to zero, a condition challenging in unreliable networks. Moreover, the nonlinearity and nonconvexity of neural networks, especially NeRF, preclude guaranteed global solutions and linear rates [10, 35, 36]. Despite these issues, Alg. 1 has proven effective in practice for distributed NeRF training, as shown in Sec. IV, converging to solutions equivalent to when a robot is doing all processing.
@@ IV Experiments
This section presents the experimental results on synthetic and real-world datasets, including 1) comparing Di-NeRF and a baseline, 2) examining the impact of the number of robots and communication graphs, 3) analyzing the impact of the amount of overlap between the robots’ trajectories on convergence, 4) evaluation using Waymo Block-NeRF dataset [8], 5) convergence of the relative pose with respect to different initial values and ground truth relative pose, 6) examining no overlap setup, 7) assessing reference robot failure, and 8) an ablation study. The architecture of NeRF is based on the instant neural graphics primitives (iNGP) [18], described in Sec IV-H. For all comparisons, the ‘baseline’ is training one model with all images, with known poses, equivalent to the standard NeRF training of iNGP. For the relative poses, the initial values for both the translational and rotational components were set to zero, except for convergence analyses (Sec. IV-E and Sec. IV-I, Fig. 14). The ground truth relative poses assume each robot’s local origin is its first frame, chosen for the largest relative pose. In reality, any image in the robot’s trajectory could serve as the origin.
IV-A Di-NeRF vs Baseline NeRF on Synthetic Dataset
@@
In this experiment, we use Di-NeRF to learn a model in a distributed manner and compare the results with the baseline, iNGP [18]. In this experiment, for Di-NeRF, the robots do not have access to a global coordinate frame, so the relative poses are unknown. We also compare Di-NeRF with PixelNeRF [16]. PixelNeRF works with one/fewer images. So part of the data from both robots that produce the best results is considered for PixelNeRF. Here, the Chair, Hotdog, and Lego sequences from the synthetic NeRF dataset [2] are used. We also use the long sequence of Barn from the real-world Tanks and Temples Benchmark [37]. The configurations for all experiments and the baseline are the same, described in Sec. IV-H. The baseline contains 100-150 images for synthetic sequences and 300 for the Barn sequence. For multiple robots, images are divided with controlled overlaps. All experiments follow official pre-processing and train/test splits.
The metrics used to compare the performance of Di-NeRF with the baseline solution are the average rendering i) PSNRs, ii) SSIM, and iii) relative pose accuracy (in degrees and cm). The pose accuracy is done against the ground truth. For the synthetic sequences, two robots split the dataset, with robot $R^{1}$ observing the chair’s front and robot $R^{2}$ the back, without trajectory overlap. This pattern applies to other sequences. The dataset is divided, and COLMAP estimates poses for each segment, with different coordinate systems. Di-NeRF calculates the relative pose. Fig. 4 shows the convergence of the relative pose for the chair sequence.
@@
Figure 3: Optimizing the relative poses for two robots for the chair sequence [2]. The relative translation and rotation estimates are plotted. The desired translation/rotation is marked by a dashed line.
@@
Figure 4: Two robots learn the scene collaboratively where each robot only sees a part of the scene. Individual training results (using iNGP and PixelNeRF) have poor quality for some areas, but Di-NeRF maintains good quality, similar to the baseline. In each column, the left and right images are for robots $R^{1}$ and $R^{2}$, respectively.
@@
Table I: Image and relative camera pose metric performances for two robots. PSNR and SSIM are reported for each robot. The relative pose accuracy is shown for $R^{2}$. $R^{1}$ is the global frame. lower/higher better is indicated by $\downarrow$/$\uparrow$ respectively.
Dataset $T_{i}^{g}$ Di-NeRF Centralized Baseline PixelNeRF
@@
Chair Known 31.93/32.05 0.941/0.945 NA NA 33.02 0.952 NA NA 33.02 0.952 24.61 0.682
Unknown 32.11/31.89 0.933/0.939 1.2 2.02 32.63 0.923 0.11 0.86
Hotdog Known 31.02/31.10 0.951/0.952 NA NA 32.63 0.963 NA NA 32.63 0.963 24.11 0.695
Unknown 30.25/30.50 0.931/0.917 1.30 1.78 32.21 0.958 0.54 0.81
Lego Known 29.00/28.91 0.936/0.937 NA NA 29.67 0.941 NA NA 29.67 0.941 25.07 0.706
Unknown 29.11/29.19 0.932/0.938 1.42 1.22 29.71 0.943 0.51 0.76
Barn Known 29.01/28.79 0.884/0.886 NA NA 29.94 0.894 NA NA 29.94 0.894 24.16 0.659
Unknown 29.09/29.23 0.854/0.843 1.42 1.70 29.66 0.891 0.59 0.94
Figure 5: NeRF validation Loss and PSNR for the chair sequence.
Fig. 4 shows the qualitative results. It illustrates the Chair sequence, comparing individual training, PixelNeRF, the baseline, and Di-NeRF for two robots. Since robot $R^{1}$ has no image from the back of the chair, it fails to reconstruct the back. Similarly, robot $R^{2}$ is unable to accurately reconstruct the front of the chair. This issue is evident in the first column of Fig. 4. The second column is PixelNeRF. The third column shows the baseline. Di-NeRF enables all robots to render the whole scene, see Fig. 4-(last column). Fig. 5 presents a comparison of PSNR and L2 loss convergences for Di-NeRF and the baseline training. The convergences exhibit nearly identical behavior.
Table I shows the quantitative results for PixelNeRF, a centralized method, the baseline, and Di-NeRF. The metrics separated by the / sign show the metric values for each robot. The centralized column refers to sending all local images and local poses to a server for training a NeRF model. The relative poses are calculated from RootSIFT features of a set of keyframes from each local set. The global poses then are calculated to train a model. Di-NeRF achieves similar quality compared to the baseline and centralized methods. For iNGP and PixelNeRF, image poses are in the global frame. Table I shows that while PixelNeRF is designed for a few input images, Di-NeRF outperforms it with results matching centralized and baseline approaches.
@@
Figure 6: (a) sample images, from the Barn sequence, provided to each robot; (b, c, d) sample images generated by Di-NeRF, once the robots collaboratively learn the scene. The images are strategically chosen—each is only visible in the raw data of a specific robot but can be rendered by all robots.
@@
Figure 7: Di-NeRF for different numbers of robots (all fully connected); (left) the allocation of frames and poses to the robots; (right) none of the robots have common images, but all robots can render the whole scene.
IV-B Di-NeRF vs Baseline NeRF on Real-world Dataset
This experiment shows the performance of Di-NeRF not only on real-world data, but also when the number of robots increases from two to five. This scenario applies to large-scale environments. Here, the Barn sequence [37] is used. There are 300 images in this sequence. Images are divided by the number of robots such that there is no common image in the local data of each robot, and the robots’ trajectories have zero overlaps. Fig. 7 shows the rendered images from each robot’s distributedly trained NeRF. Each robot can render a view from any point (e.g. Robot $R^{3}$ has never seen the large bay doors but renders an image when requested, as seen in Fig. 7 (b)).
Next, the number of robots in this experiment is altered. In Fig. 7, the trajectory for each robot is shown alongside the rendered images for each robot after collaborative training is completed. Each color represents the trajectory of a different robot. Table II presents the quantitative results. The rendering qualities and accuracy of relative camera pose are compared to the baseline method. With Di-NeRF, the outcomes closely resemble those achieved with the baseline method. In Table III, we experiment with different connectivity graphs of robots. Five robots are used with the Barn sequence. In the star connection, $R^{1}$ is in the center. In Table III, the average results for 5 robots are shown.
Table II: Image and relative camera pose metric performances for different numbers of robots for the Barn sequence.
Di-NeRF Baseline Di-NeRF Baseline Relative Camera Pose Error
#Robots PSNR $\uparrow$ SSIM $\uparrow$ $\delta R^{\circ}$downarrow $\delta t(cm)$downarrow
2 29.21/29.11 29.94 0.886/0.839 0.894 1.34 1.76
3 29.13/29.21/29.08 0.836/0.878/0.873 2.12/1.89 1.05/1.82
4 29.35/29.39/29.33/29.17 0.856/0.838/0.828/0.856 1.35/1.90/2.10 1.83/2.00/3.01
5 29.67/29.44/29.15/29.45/29.41 0.845/0.878/0.852/0.887/0.869 2.70/3.20/3.10/3.00 2.11/2.65/2.39/2.02
Table III: Mean validation result of Di-NeRF for 5 robots, with different connectivities, for the Barn sequence. Di-NeRF works well even if the network of robots is not fully connected. lower/higher better is indicated by $\downarrow$/$\uparrow$ respectively.
[…]
Fully Connected 29.42 0.866 3.00 2.29
Ring 29.19 0.835 4.56 2.33
Star 29.16 0.862 4.91 3.02
Line 29.10 0.853 5.01 2.97
Figure 8: Segmenting the Chair sequence with different overlaps.
Figure 9: Optimizing the relative poses for two robots with varying overlaps.